ABSOLUTE LINE INTENSITIES FOR THE ν_{6} BAND OF H_{2}O_{2}

S. KLEE, M. WINNEWISSER, Justus Liebig Universita̋t, , Physikalisch-Chemisches Institut Heinrich-Buff-Ring 58, D-35392 Gießen, Deutschland; \underline{A. PERRIN}, J.-M. FLAUD, Laboratoire de Photophysique Moléculaire, CNRS, Université Paris Sud, Campus d’Orsay, Bat 210, 91405 Orsay Cedex, France.

The purpose of this work was to obtain reliable absolute intensities for the ν_{6} band of H_{2}O_{2}. This work was undertaken because strong discrepancies exist between the different ν_{6} band intensities which are presently available in the literature ^{a} ^{b}. The method which was used in the present work is to use Fourier transform spectra of H_{2}O_{2} recorded at Giessen in a spectral range (370-1270cm^{−1}) which covers both the R-branch of the torsion-rotation band at low wavenumber and the P-branch of the ν_{6} band at high wavenumber.

• From the low wavenumber data, the partial pressure of H_{2}O_{2} present in the cell during the recording of the spectra was determined by calibrating the observed absorptions in the torsion-rotation band with intensities computed ^{c} by using the permanent H_{2}O_{2} dipole moment measured by Stark effect ^{d}.

• In the high frequency range, this value of the partial pressure of H_{2}O_{2} was used to measure absolute line intensities in the ν_{6} band.

Finally, the line intensities in the ν_{6} band were fitted using the theoretical methods described in detail in our previous works ^{a,c}.

aA. Perrin, A. Valentin, J.-M. Flaud, C. Camy-Peyret, L. Schriver, A. Schriver, and Ph. Arcas, J. Mol. Spectrosc. \underline{171}, 358 (1995).

bR. May, J. Quant. Radiat. Transfer. \underline{45}, 267 (1991).

cA. Perrin, J.-M. Flaud, C. Camy-Peyret, R. Schermaul, M. Winnewisser, J. Y. Mandin, V. Dana, M. Badaoui, and J. Koput, J. Mol. Spectrosc. \underline{176}, 287 (1996).

dE.A. Cohen, and H.M. Pickett, J. Mol. Spectrosc. \underline{87}, 582 (1981).